OpenStudy (anonymous):

the problem is:the perimeter of a rectangle is 408cm and the length is 134cm. what is the width and area

OpenStudy (anonymous):

 now how many sides are there in the rectangle

OpenStudy (anonymous):

4

OpenStudy (anonymous):

k h0w many L and how many W

OpenStudy (anonymous):

i think i had to write P=2(L+w)

OpenStudy (anonymous):

gr8

OpenStudy (anonymous):

so would p=2(408+x) 2x134+x=x

OpenStudy (anonymous):

no u r doing wrong

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Here, perimeter= 408cm and the length = 134cm

OpenStudy (anonymous):

yeah i know that part p=408 l=134 w=x

OpenStudy (anonymous):

thats wat my teacher told me. i dont know what too do after that

OpenStudy (anonymous):

did u got the answer. in which class do u study

OpenStudy (anonymous):

algabra 1

OpenStudy (anonymous):

how would you put that as an equation